OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 840! If three times the number of one-step equations minus twice the number of two-step equations is equal to 870, how many one-step equations auditioned today?

hartnn (hartnn):

hint,let number of 1 step eq be a and number of 2 step eq be b so,a+b=840 3a-2b=870 can u solve these 2 equations to get a and b?

hero (hero):

If I had a nickel for every time I saw this question.

OpenStudy (anonymous):

i don't get how to solve the equations

OpenStudy (anonymous):

@hartnn how would i solve them?

hartnn (hartnn):

u can first find a from a+b=840 so tell me what u get as a?

OpenStudy (anonymous):

a=840-b?

hartnn (hartnn):

yes :) now put it in 2nd equation 3(840-2b)-2b=870 can u get b?

OpenStudy (anonymous):

b=-206.25

hartnn (hartnn):

isn't it +206.25

OpenStudy (anonymous):

yes it is i found the mistake

hartnn (hartnn):

now just put the value of b in a+b=840 to get a

OpenStudy (anonymous):

i got 633.75 but my only choices are 330, 510, 520, 660

OpenStudy (anonymous):

i got the answer thank you